Two sides of a triangle have the same length. the third side measures 5m less than twice the common length. The perimeter is 11m. What is the length of the three sides?
Answer by nerdybill(7384)

Two sides of a triangle have the same length. the third side measures 5m less than twice the common length. The perimeter is 11m. What is the length of the three sides?
.
Let x = measure of sides that have common length
then from "the third side measures 5m less than twice the common length."
2x-5 = third side
.
Because they gave us the perimeter we have:
11 = x + x + 2x-5
11 = 4x-5
16 = 4x
4 m = x
.
third side:
2x-5 = 2(4)-5 = 8-5 = 3 m
.
Lengths of the three sides:
4m, 4m, and 3m